OpenStudy (anonymous):

Verify identity cot^2x + csc^2x = 2csc^2x - 1

OpenStudy (lgbasallote):

\[\cot^2 x + \csc^2 x\] \[\implies \frac{\cos^2 x} {\sin^2 x} + \frac{1}{\sin^2 x}\] \[\implies \frac{\cos^2 x +1}{\sin^2 x}\] \[\implies \frac{1- \sin^2 x + 1}{\sin^2 x}\] \[\implies \frac{2 - \sin^2 x}{\sin^2 x}\] \[\implies \frac{2}{\sin^2 x} - \frac{\sin^2 x}{\sin^2 x}\] \[\implies 2 \csc^2 x - 1\] does that help?

OpenStudy (anonymous):

thank u!

OpenStudy (lgbasallote):

welcome :D